Solve.
4-sqrt(4+x^2)=x
Thank you in advance for any help:)

Answer by Boreal(15235)

4-sqrt(4+x^2)=x
4-x=sqrt(4+x^2)
square both sides
16-8x+x^2=4+x^2
12=8x
x=3/2 or 1.5
check
4-sqrt(6.25)=1.5
sqrt(6.25)=2.5
4-2.5=1.5 check